OpenStudy (anonymous):

father age is 3 times as his son age. after 15 years fathers age is double of son age. find the current ages of father and son.

OpenStudy (anonymous):

Father's age = y, sons age = x. y=3x and y=2x+15. Find the intersection: 3x = 2x +15, x =15. The son is 15. Substitute back into y=3x and the father is 45. Check: y=2x + 15, 45=2(15)+15 = 45. Correct.

OpenStudy (anonymous):

3x = y y + 15 = 2(x + 15) 3x + 15 = 2x +30 x = 15 y = 45

OpenStudy (anonymous):

45 age for father and 15 for son